# Floquet breathers in a modulated nonlinear lattice

### v3

@@ Appendix A Numerical experiments, thermalization and noisy systems
We perform numerical simulations and analytical calculations to obtain properties of the system as the initial coordinates and momenta for breathers, or the discrete Fourier components, or to find their stability properties. These simulations are supposed to produce a definite and precise outcome. The imprecission is due to shortcomings of the numerical method, the parameters chosen, as time integration steps, the number of Fourier components and so on.
@@
Therefore, we consider that the system is in thermal equilibrium when some of its consequences are approximately met. We understand that (i) the system forgets its microscopic initial conditions, i.e., positions and velocities but not energy and other quantities; (ii) its local energy probability distribution approaches the Boltzmann distribution; (iii) identical particles have the same properties over time; (iv) the participation function aproaches $N/2$ (see below); (v) the site of the particle with peak local energy moves along the lattice, perhaps with space-time correlation of some time when fluctuations bring about wave groups with some lifetime (See. Fig. 2) but not permanently unless they are a consequence of the system properties. Some temporary correlations are also properties of the system as phonon wave groups. Those are detected to construct, for example, the numerically experimental dispersion relations as explained in App. C.1.
This subject is dealt in Ref. Archilla et al., 2025 in the following way. Let us consider a system of particles with weak coupling in the microcanonical ensemble, we can suppose that each particle is in contact with a thermal bath corresponding to the rest of the system at a given temperature $T$, or thermodynamic $\beta=1/\mu$, with $\mu=kT$, $k$ being the Boltzmann constant. Within this approximation, the probability of a particle having an energy $e_{n}$ is proportional to $\exp(-\beta e_{n})$,Reif (2009) with $\mu$ equal to twice the average kinetic energy, according to the virial theorem.
@@
Imposing the probability one for a particles to have all the energies from 0 to $\infty$, the probability density of a particle having energy $e_{n}$ is $\rho(e_{n})=\exp(-e_{n}/\mu)/\mu$, also known as the exponential distribution. The average particle energy is $\langle e_{n}\rangle=\mu$, with standard deviation also $\mu$. Averages are in principle for the statistical ensemble, but due to the ergodic theorem, the average can be made in time and also at the different identical particles.
Figure 16: (Top:) Path to thermal equilibrium for the particles outside a breather core. Initial conditions are a random distribution of velocities with the coordinates at rest except for the initial breather coordinates at a core of $1$ particles in a system with $N=160$ particles. The theoretical value at equilibrium is $P=N/2$ (in red), but with large fluctuations for a N=16011=149N=160-11=149 system. In this case, a shorter time than in Fig. 15 is shown in linear scale. (Bottom:) Numerical experimental distribution function (blue) compared with the theoretical one (red) outside the breather core. Parameters: $N=160$, $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\Omega=\omega_{\text{b}}/20=0.049$, $h=0$.
An easily measurable consequence of this distribution is the behavior of the participation function $P=(\sum e_{n})^{2}/\sum e_{n}^{2}$. This is a function with value 1 if all the energy is located at a single particle and $N$ if the energy is evenly distributed among the system number of particles $N$. The value of $P$ at thermal equilibrium using the exponential distribution is $N/2$, with relatively large oscillations of $P$ due to the large fluctuations of $e_{n}$ for the exponential distribution. The amplitude of the oscillation of $P$ around $N/2$ depends on the system size, the larger the size, the relative value of the oscillation becomes smaller as the relative standard deviation is or the order of $1/\sqrt{N}$. As the distribution $\rho(e_{n})$ is not exact, the actual value of the average $P$ can be a few particles above or below $N/2$, also changing with the time observation window, due to the large fluctuations of a small system with $N\simeq 100$.
We can check these properties for our system with different types of modulation. The exponential distribution fails at low energies when the model of a particle with energy $e_{n}$ interacting with the nearest neighbors is not correct as part of the energy is precisely the interaction energy with the nearest neighbors. Often the first bin to obtain the numerical distribution have to be discarded.
@@
The concept of noisy system appears when we study breathers. The linear stability of breathers, and its nonlinear stability for small perturbations, is analyzed in Sec. IV.D, but another complementary technique of particular physical interest is to observe the breather behavior in presence of temperature. A system with a breather is not thermalized, because eventually the breather will disappear but the breather life can be very long for some systems and breather energies.Archilla et al. (2025) A noisy system is a system that it is in an approximate thermal equilibrium, in the sense explained above, outside the breather core. The breather core is a few particles around the breather center, depending on the breather localization.
@@ Appendix B Notation
To present formally the theory, we will use the ket and bra notation. The nice thing of that notation is that the properties of operations with vectors, change of basis, and operators appear as the simple juxtaposition of kets and bras. We present here this notation, more details can be read in Ref. Griffiths, 2024 and in Ref. Archilla et al., 2003 applied to breather theory.
@@
The bra $\langle\phi|$ is the Hermitian conjugate, so $\langle\phi|$ is a row matrix with elements the complex conjugates of the elements of $|\phi\rangle$. In this way, a bracket $\langle\phi_{1}|\phi_{2}\rangle$ is the scalar product $\sum_{n=0}^{N-1}\phi_{1,n}^{*}\phi_{2,n}$. Trivially $\langle n|n^{\prime}\rangle=\delta_{n,n^{\prime}}$. So the basis $\{|n\rangle\}$ is orthonormal. $\langle n|\phi\rangle=\phi_{n}$ is the scalar product by a unit basis vector and therefore the component of $|\phi\rangle$ on that basis vector.
@@ Appendix C Derivation of the effect of space-time modulation on the linear dispersion
The dispersion relations obtained in numerical experiments and the theoretical ones have been already presented in Sec. III, with the exception of only space modulation presented below as it was not essential for the rest of the article. Here, after some short introduction, we describe the method to visualize the dispersion relations in numerical experiments followed by the theoretical derivations in several cases.
@@ C.1 Obtention of the dispersion curves in numerical experiments
We obtain numerically what we could call the experimental dispersion curves in an approximate thermal equilibrium, as explained with detail in App. A, with the following procedure: First we consider a relatively large system, with, for example, $N=128$ oscillators, and use as initial conditions a random distribution of positions for the variables $u_{n}$ in [-0.5,0.5], Equally, the random perturbation can be added to the momenta or both the momenta and coordinates without changing the outcome. Second, we leave the linear system (7) to evolve a relatively long time, $t=5000$, for example, when the system is supposed to be near or at thermal equilibrium and therefore all the linear modes or phonons are in principle excited. We can use also the full nonlinear system but with conveniently small values of the coordinates or momenta so that the system is near the linear limit. Then, we use the final positions and velocities, to run a simulation of similar length. We use the results to perform the two dimensional discrete Fourier transform in space and time (XTDFT) on the variable $u_{n}(t)$, that is,
@@
with $t_{l}=\frac{\displaystyle l}{\displaystyle N_{t}}t_{f}$, $\omega_{r}=\frac{\displaystyle r}{\displaystyle N_{t}}\frac{\displaystyle 2\pi}{\displaystyle\Delta t}$, for $l,r=0\dots N_{t}-1$ and $q_{k}=\frac{\displaystyle k}{\displaystyle N}2\pi$, for $k=0\dots N-1$. The sampling interval $\Delta t=\frac{\displaystyle t_{f}}{\displaystyle N_{t}}$ is larger that the integration step to avoid very large matrices as the maximum frequency $2\pi/\Delta t$ does not need to be very large because the frequencies are of the order of $\omega_{0}$ and $\Omega$. On the other hand the frequency precision $\Delta\omega=\frac{\displaystyle 2\pi}{\displaystyle N_{t}\Delta t}=\frac{\displaystyle 2\pi}{\displaystyle t_{f}}$, is inversely proportional to $t_{f}$, therefore a large value of it is convenient, both to obtain precision for the frequency and for the results to be generic enough.
The intensities are defined as $I(\omega_{i},q_{n})=|F(\omega_{i},q_{n})|^{2}$, normalized to unity. To visualize the dispersion curves, we represent a number of lines in contour plots, typically ten, to get a clear picture. However, all the modes do not have the same probability to occur, the density of states may be small, and theoretically existing modes may have small intensities and are difficult to discriminate from the noise background, those dispersion lines are weak. Presently, we do not have a means to calculate theoretically those intensities.
@@ C.2 Dispersion bands modification with only space modulation
In order to understand the result we will consider first the case $h=2\pi/L$, where $L$ is an integer, and $\Omega=0$. In this case, the symmetry of the system $n\rightarrow n+1$ is broken and substituted by the symmetry $n\rightarrow n+L$. Let us suppose that $\delta=0$, but still the symmetry is broken, then, in order to be able to impose periodic boundary conditions, we need that $N=RL$ with $R$ and integer number. Now, by Bloch theorem,Griffiths and Schroeter (2018) the solutions are of the form $\exp(\textrm{i}qn)F(n,t)$, with $F(n,t)$ a solution with the periodicity of the lattice, whose period is now $L$, that is the elements of the basis are given by $u_{l,r}=\exp(\textrm{i}q_{r}n)\exp[\textrm{i}(q_{l}n-\omega t])$. As the second factor has the periodicity of the new lattice $q_{l}=(2\pi/L)l$, with $l=0,\dots,L-1$. If we impose periodic boundary conditions, then $u_{l,r}(n,t)=u_{l,r}(n+N,t)$, the second factor is automatically identical if $N=RL$ and the first factor becomes $q_{r}=2\pi r/N$, with $r=0,\dots,R-1$, as larger values of $r$ can be written as $q_{r}^{\prime}=2\pi(r_{1}R+r_{2})/N=2\pi(r_{1}R+r_{2})/LR=2\pi r_{1}/L+2\pi r_{2}/N$, and the first term is periodic in $L$ and can be incorporated into $F(n,t)$. We can rewrite it as:
@@
Figure 17: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for explanation. Parameters: $\kappa=0.10$, $\delta=0.05$, $\Omega=0$ .
@@ C.3 Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. They are obtained numerically as explained in App. C.1. Some bands are weak, meaning that their intensity in the XTDFT is small, presumably because their density of states is small. Some becomes so weak that they can not be seen into the background noise. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to a divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
@@
where $\bar{\omega}_{q}$ is the unperturbed frequency for the wavenumber $q$. This is a complicated equation, but if $\delta$ is small with respect to $\omega_{0}`2$, we can suppress that term in the first approximation, that is.

### v4

@@ Appendix A Numerical experiments, thermalization and noisy systems
We perform numerical simulations and analytical calculations to obtain properties of the system as the initial coordinates and momenta for breathers, or the discrete Fourier components, or to find their stability properties. These simulations are supposed to produce a definite and precise outcome. The imprecision is due to shortcomings of the numerical method, the parameters chosen, as time integration steps, the number of Fourier components, and so on.
@@
Therefore, we consider that the system is in thermal equilibrium when some of its consequences are approximately met. We understand that (i) the system forgets its microscopic initial conditions, i.e., positions and velocities but not energy and other quantities; (ii) its local energy probability distribution approaches the Boltzmann distribution; (iii) identical particles have the same properties over time; (iv) the participation function approaches $N/2$ (see below); (v) the site of the particle with peak local energy moves along the lattice, perhaps with space-time correlation of some time when fluctuations bring about wave groups with some lifetime (See. Fig. 2) but not permanently unless they are a consequence of the system properties. Some temporary correlations are also properties of the system as phonon wave groups. Those are detected to construct, for example, the numerically experimental dispersion relations as explained in App. C.1.
This subject is dealt with in Ref. Archilla et al., 2025 in the following way. Let us consider a system of particles with weak coupling in the microcanonical ensemble, we can suppose that each particle is in contact with a thermal bath corresponding to the rest of the system at a given temperature $T$, or thermodynamic $\beta=1/\mu$, with $\mu=kT$, $k$ being the Boltzmann constant. Within this approximation, the probability of a particle having an energy $e_{n}$ is proportional to $\exp(-\beta e_{n})$,Reif (2009) with $\mu$ equal to twice the average kinetic energy, according to the virial theorem.
@@
Imposing the probability one for a particle to have all the energies from 0 to $\infty$, the probability density of a particle having energy $e_{n}$ is $\rho(e_{n})=\exp(-e_{n}/\mu)/\mu$, also known as the exponential distribution. The average particle energy is $\langle e_{n}\rangle=\mu$, with standard deviation also $\mu$. Averages are in principle for the statistical ensemble, but due to the ergodic theorem, the average can be made in time and also at the different identical particles.
Figure 16: (Top:) Path to thermal equilibrium for the particles outside a breather core. Initial conditions are a random distribution of velocities with the coordinates at rest except for the initial breather coordinates at a core of $1$ particles in a system with $N=160$ particles. The theoretical value at equilibrium is $P=N/2$ (in red), but with large fluctuations for a N=16011=149N=160-11=149 system. In this case, a shorter time than in Fig. 15 is shown on a linear scale. (Bottom:) Numerical experimental distribution function (blue) compared with the theoretical one (red) outside the breather core. Parameters: $N=160$, $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\Omega=\omega_{\text{b}}/20=0.049$, $h=0$.
An easily measurable consequence of this distribution is the behavior of the participation function $P=(\sum e_{n})^{2}/\sum e_{n}^{2}$. This is a function with value 1 if all the energy is located at a single particle and $N$ if the energy is evenly distributed among the system number of particles $N$. The value of $P$ at thermal equilibrium using the exponential distribution is $N/2$, with large oscillations of $P$ due to the large fluctuations of $e_{n}$ for the exponential distribution. The amplitude of the oscillation of $P$ around $N/2$ depends on the system size, the larger the size, the relative value of the oscillation becomes smaller as it is of the order of $1/\sqrt{N}$. As the distribution $\rho(e_{n})$ is not exact, the actual value of the average $P$ can be a few particles above or below $N/2$, also changing with the time observation window, due to the large fluctuations of a small system with $N\simeq 100$.
We can check these properties for our system with different types of modulation. The exponential distribution fails at low energies when the model of a particle with energy $e_{n}$ interacting with the nearest neighbors is not correct as part of the energy is precisely the interaction energy with the nearest neighbors. Often the first bin to obtain the numerical distribution has to be discarded.
@@
The concept of a noisy system appears when we study breathers. The linear stability of breathers, and its nonlinear stability for small perturbations, is analyzed in Sec. IV.D, but another complementary technique of particular physical interest is to observe the breather behavior in presence of temperature. A system with a breather is not thermalized, because eventually the breather will disappear but the breather life can be very long for some systems and breather energies.Archilla et al. (2025) A noisy system is a system that it is in an approximate thermal equilibrium, in the sense explained above, outside the breather core. The breather core is a few particles around the breather center, depending on the breather localization.
@@ Appendix B Notation
To present formally the theory, we will use the ket and bra notation. The nice thing of that notation is that the properties of operations with vectors, change of basis, and operators appear as the simple juxtaposition of kets and bras. We present here this notation; more details can be read in Ref. Griffiths, 2024 and in Ref. Archilla et al., 2003 applied to breather theory.
@@
The bra $\langle\phi|$ is the Hermitian conjugate, so $\langle\phi|$ is a row matrix with elements the complex conjugates of the elements of $|\phi\rangle$. In this way, a bracket $\langle\phi_{1}|\phi_{2}\rangle$ is the scalar product $\sum_{n=0}^{N-1}\phi_{1,n}^{*}\phi_{2,n}$. Trivially $\langle n|n^{\prime}\rangle=\delta_{n,n^{\prime}}$. So the basis $\{|n\rangle\}$ is orthonormal. $\langle n|\phi\rangle=\phi_{n}$ is the scalar product by an unit basis vector and, therefore, the component of $|\phi\rangle$ on that basis vector.
@@ Appendix C Derivation of the effect of space-time modulation on the linear dispersion
The dispersion relations obtained in numerical experiments and the theoretical ones have already been presented in Sec. III, with the exception of only space modulation presented below as it was not essential for the rest of the article. Here, after some short introduction, we describe the method to visualize the dispersion relations in numerical experiments followed by the theoretical derivations in several cases.
@@ C.1 Obtention of the dispersion curves in numerical experiments
We obtain numerically what we could call the experimental dispersion curves in an approximate thermal equilibrium, as explained with detail in App. A, with the following procedure: First we consider a relatively large system, with, for example, $N=128$ oscillators, and use as initial conditions a random distribution of positions for the variables $u_{n}$ in [-0.5,0.5], Equally, the random perturbation can be added to the momenta or both the momenta and coordinates without changing the outcome. Second, we leave the linear system (7) to evolve a relatively long time, $t=5000$, for example, when the system is supposed to be near or at thermal equilibrium and, therefore, all the linear modes or phonons are in principle excited. We can also use the full nonlinear system but with conveniently small values of the coordinates or momenta so that the system is near the linear limit. Then, we use the final positions and velocities, to run a simulation of similar length. We use the results to perform the two-dimensional discrete Fourier transform in space and time (XTDFT) on the variable $u_{n}(t)$, that is,
@@
with $t_{l}=\frac{\displaystyle l}{\displaystyle N_{t}}t_{f}$, $\omega_{r}=\frac{\displaystyle r}{\displaystyle N_{t}}\frac{\displaystyle 2\pi}{\displaystyle\Delta t}$, for $l,r=0\dots N_{t}-1$ and $q_{k}=\frac{\displaystyle k}{\displaystyle N}2\pi$, for $k=0\dots N-1$. The sampling interval $\Delta t=\frac{\displaystyle t_{f}}{\displaystyle N_{t}}$ is larger than the integration step to avoid very large matrices as the maximum frequency $2\pi/\Delta t$ does not need to be very large because the frequencies are of the order of $\omega_{0}$ and $\Omega$. On the other hand the frequency precision $\Delta\omega=\frac{\displaystyle 2\pi}{\displaystyle N_{t}\Delta t}=\frac{\displaystyle 2\pi}{\displaystyle t_{f}}$, is inversely proportional to $t_{f}$, therefore, a large value of it is convenient, both to obtain precision for the frequency and for the results to be generic enough.
The intensities are defined as $I(\omega_{i},q_{n})=|F(\omega_{i},q_{n})|^{2}$, normalized to unity. To visualize the dispersion curves, we represent a number of lines in contour plots, typically ten, to get a clear picture. However, all the modes do not have the same probability to occur, the density of states may be small, and theoretically existing modes may have small intensities and are difficult to discriminate from the noise background, those dispersion lines are weak. Presently, we do not have a means to calculate those intensities theoretically.
@@ C.2 Dispersion bands modification with only space modulation
In order to understand the results, we will consider first the case $h=2\pi/L$, where $L$ is an integer, and $\Omega=0$. In this case, the symmetry of the system $n\rightarrow n+1$ is broken and substituted by the symmetry $n\rightarrow n+L$. Let us suppose that $\delta=0$, but still the symmetry is broken, then, in order to be able to impose periodic boundary conditions, we need that $N=RL$ with $R$ and integer number. Now, by Bloch theorem,Griffiths and Schroeter (2018) the solutions are of the form $\exp(\textrm{i}qn)F(n,t)$, with $F(n,t)$ a solution with the periodicity of the lattice, whose period is now $L$, that is the elements of the basis are given by $u_{l,r}=\exp(\textrm{i}q_{r}n)\exp[\textrm{i}(q_{l}n-\omega t])$. As the second factor has the periodicity of the new lattice $q_{l}=(2\pi/L)l$, with $l=0,\dots,L-1$. If we impose periodic boundary conditions, then $u_{l,r}(n,t)=u_{l,r}(n+N,t)$, the second factor is automatically identical if $N=RL$ and the first factor becomes $q_{r}=2\pi r/N$, with $r=0,\dots,R-1$, as larger values of $r$ can be written as $q_{r}^{\prime}=2\pi(r_{1}R+r_{2})/N=2\pi(r_{1}R+r_{2})/LR=2\pi r_{1}/L+2\pi r_{2}/N$, and the first term is periodic in $L$ and can be incorporated into $F(n,t)$. We can rewrite it as:
@@
Figure 17: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for the explanation. Parameters: $\kappa=0.10$, $\delta=0.05$, $\Omega=0$ .
@@ C.3 Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. They are obtained numerically as explained in App. C.1. Some bands are weak, meaning that their intensity in the XTDFT is small, presumably because their density of states is small. Some become so weak that they cannot be seen into the background noise. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to the divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
@@
where $\bar{\omega}_{q}$ is the unperturbed frequency for the wavenumber $q$. This is a complicated equation, but if $\delta$ is small with respect to $\omega_{0}^{2}$, we can suppress that term in the first approximation, that is.